# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 2 Low-rank approximations to linear Gaussian inverse problems
Let $\mathcal{H}$ be a separable Hilbert space over $\mathbb{R}$ of dimension $\dim{H}\leq\infty$. Consider the observation model defined by a continuous linear forward model $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ and additive Gaussian observation error
@@
The covariance $\mathcal{C}_{\textup{obs}}\in B(\mathbb{R}^{n})_{\mathbb{R}}$ of the observation noise $\zeta$ is positive, and from a frequentist nonparametric perspective, $x^{\dagger}\in\mathcal{H}$ is the unknown true data-generating parameter to recover after observing a realisation of $Y$. By the Gaussian assumption on the noise, it follows that for any fixed $x\in\mathcal{H}$, the likelihood of observing $y$ is proportional to $\exp(-\frac{1}{2}\lVert\mathcal{C}_{\textup{obs}}^{-1}(y-Gx)\rVert^{2})$. The Hessian of the negative log-likelihood is
@@
We adopt the Bayesian perspective to the problem of inferring $x^{\dagger}$ given the observation $y$ of $Y$, by modeling the unknown $x^{\dagger}$ with an $\mathcal{H}$-valued random variable $X$. Its distribution, the prior distribution, is taken to be a nondegenerate Gaussian measure $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$ on $\mathcal{H}$ and we assume that $X$ and $\zeta$ are independent. As the covariance of a Gaussian measure, $\mathcal{C}_{\textup{pr}}$ lies in $L_{1}(\mathcal{H})_{\mathbb{R}}$ and $\mathcal{C}_{\textup{pr}}\geq 0$, hence $\mathcal{C}_{\textup{pr}}$ has a unique nonnegative square root $\mathcal{C}_{\textup{pr}}^{1/2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$. Nondegeneracy of $\mu_{\textup{pr}}$ implies that $\operatorname{supp}(\mu_{\textup{pr}})=\mathcal{H}$, see e.g. [4, Definition 3.6.2] and that $\mathcal{C}_{\textup{pr}}$ is injective, hence $\mathcal{C}_{\textup{pr}}>0$ and $\mathcal{C}_{\textup{pr}}^{1/2}>0$, see Lemma A.19. Hence the inverse $\mathcal{C}_{\textup{pr}}^{-1}$ is a well-defined bijection $\operatorname{ran}{C}_{\textup{pr}}\rightarrow\mathcal{H}$ which is self-adjoint, c.f. Definition A.12 and Lemma A.16(ii), and is unbounded if $\dim{H}=\infty$. The Cameron–Martin space of $\mu_{\textup{pr}}$ is the Hilbert space $(\operatorname{ran}{C}_{\textup{pr}}^{1/2},\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}})$, see e.g. [4, p. 293], where the Cameron–Martin norm of an element $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ is defined by $\lVert h\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}\coloneqq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert$. As $\mathcal{C}_{\textup{pr}}$ is injective and compact, $\operatorname{ran}{C}_{\textup{pr}}^{1/2}$ is dense in $\mathcal{H}$ and if $\dim{H}=\infty$, then $\operatorname{ran}{C}_{\textup{pr}}^{1/2}$ is a proper dense subspace of $\mathcal{H}$.
Given a realisation $y$ of the random variable $Y$ defined by the observation model, the posterior law $\mu_{\textup{pos}}=\mu_{\textup{pos}}(y)$ of $X$ given $Y=y$ is the Gaussian measure $\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$, where
@@
see e.g. [31, Example 6.23]. Equation (3c) should be understood to hold on $\operatorname{ran}{C}_{\textup{pos}}$. While all nondegenerate Gaussians are equivalent in finite-dimensionsional setting, this is no longer true in infinite-dimenional setting, where in fact it holds that nondegenerate Gaussians that are not equivalent must be mutually singular. By [31, Section 6.3], $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are equivalent. We shall construct approximations to $\mu_{\textup{pos}}$ that are equivalent to $\mu_{\textup{pos}}$.
@@
The equations in (3) motivate certain Gaussian approximations of $\mu_{\textup{pos}}$ that, as we shall see, retain equivalence to $\mu_{\textup{pos}}$. By (3b), $\mathcal{C}_{\textup{pos}}$ is an update of $\mathcal{C}_{\textup{pr}}$ by a nonpositive self-adjoint operator $-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}}$. The range of this update is contained in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and the rank of this update is at most $n$ since $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{r})$. For $r\in\mathbb{N}$, this motivates the rank-constained approximation of $\mathcal{C}_{\textup{pos}}$ by updating $\mathcal{C}_{\textup{pr}}$ by nonpositive self-adjoint operators of the form $KK^{*}$, for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}$. That is, we consider
@@
$\mathscr{C}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\right\},\quad r\in\mathbb{N},$ (4)
and we have $\mathcal{C}_{\textup{pos}}\in\mathscr{Cj_{r}}$ for some $r\leq n$.
Alternatively, we can consider approximations of $\mu_{\textup{pos}}$ by constructing rank-constrained updates of the prior precision $\mathcal{C}_{\textup{pr}}^{-1}$. By (3c), $\mathcal{C}_{\textup{pr}}^{-1}$ is an update of $\mathcal{C}_{\textup{pr}}^{-1}$ by the Hessian $H$, which is a self-adjoint, nonnegative, and has rank at most $n$. For $r\in\mathbb{N}$, we can therefore consider the class of approximations of $\mathcal{C}_{\textup{pos}}^{-1}$ of the form $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$, for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. That is, we consider
@@
$\mathscr{P}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\right\},\quad r\in\mathbb{N},$ (5)
and hence $\mathcal{C}_{\textup{pos}}^{-1}\in\mathscr{P}_{r}$ for some $r\leq n$, by the fact that $H\in\mathcal{B}_{00,n}(\mathcal{H})$.
We note that every operator $SS^{*}$ for $S\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a positive, self-adjoint operator with rank at most $r$, and that every nonnegative operator $T\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$ can be written in this way. Therefore, we could write the above approximations as $\mathcal{C}_{\textup{pr}}-T$ or $\mathcal{C}_{\textup{pr}}^{-1}+T$ for nonnegative $T\in\mathcal{B}_{00,r}(\mathcal{H})$, such that $\mathcal{C}_{\textup{pr}}-T$ is positive and maps into $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. However, the set of nonnegative elements of $\mathcal{B}_{00,r}(\mathcal{H})$ is not convex, since rank is not preserved by convex combinations. By replacing $T$ by $SS^{*}$, we avoid formulating an optimisation problem over a nonconvex set. Indeed, $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is not only convex but is also a Banach space.
For approximations of the posterior mean, we observe from (3a) that the posterior mean is the result of left-multiplying the data $y$ by an operator. We shall consider approximations of the posterior mean that retain equivalence with the exact posterior. This motivates the following classes of operators:
$\mathscr{M}^{(1)}_{r}\coloneqq$$\{(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ B\in\mathcal{B}_{00,r}(\mathcal{H}),\ N((\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}y,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}\},$ (6a)
@@
$\mathscr{M}^{(2)}_{r}\coloneqq$$\{A\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}):\ N((\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}y,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}\}.$ (6b)
The class in (6a) takes properties of the posterior covariance that are implied by (3b) into account. In contrast, the class in (6b) ignores this structure and takes only into account that the resulting approximate posterior approximation is equivalent to the exact posterior. By (3a), it follows that there exists $r\leq n$ such that $m_{\textup{pos}}\in\mathscr{M}^{(1)}_{r}\cap\mathscr{M}^{(2)}_{r}$. In fact, if $r\geq\operatorname{rank}\left({G^{*}}\right)=\operatorname{rank}\left({G}\right)$, then $(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H})$ for every $B\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$. The classes $\mathscr{C}_{r}$, $\mathscr{P}_{r}$, $\mathscr{M_{r}}^{(1)}$ and $\mathscr{M}_{r}^{(2)}$ generalise those considered in [30]. It is a natural to search for low-rank approximations of the objects in (3), where ‘low-rank’ refers to the fact that we consider approximations in the classes $\mathscr{C}_{r}$, $\mathscr{P}_{r}$, $\mathscr{M}_{r}^{(1)}$ and $\mathscr{M}_{r}^{(2)}$ for $r<n$ respectively.
@@ 3 Equivalence and Divergences between Gaussian measures
Since our approximation problems are formulated in the context of statistical inverse problems, and since absolute continuity of the posterior with respect to the prior is important for statistical inference, we ensure that the approximate posteriors we consider are equivalent to $\mu_{\textup{pos}}$. In Section 3.1, we recall the Feldman–Hajek theorem which gives necessary and sufficient conditions for Gaussian measures to be equivalent, and apply this theorem to the setting described in Section 2. Then, in Section 3.2, we consider certain divergences between equivalent Gaussian measures, which we can use to measure approximation quality of our constructed low-rank posterior approximations.
@@ 3.1 Equivalence between Gaussian measures
Let $\mathcal{H}$ be a Hilbert space and $\mu=N(m_{1},\mathcal{C}_{1})$ and $\nu=(m_{2},\mathcal{C}_{2})$ be Gaussian measures on $\mathcal{H}$. Then $\mu$ and $\nu$ are singular or equivalent, and $\mu$ and $\nu$ are equivalent if and only if the following conditions hold:
@@
$\mathcal{E}\coloneqq\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{N}(m_{\textup{pos}},\mathcal{C})\sim\mu_{\textup{pos}}\}.$ (7)
That is, $\mathcal{E}$ contains those covariances $\mathcal{C}$ such that $\mathcal{N}(m_{\textup{pos}},\mathcal{C})$ is equivalent to $\mu_{\textup{pos}}$. Since $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are equivalent, we have $\mathcal{C}_{\textup{pr}}\in\mathcal{E}$.
@@
By [4, Lemma 6.3.1(ii)] the set of operators that satisfy property E is closed under taking inverses, adjoints and compositions. Furthermore, since $\mathcal{\mu}_{\textup{pos}}\sim\mu_{\textup{pr}}$, $\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}$ satisfies property E. One can now use Theorem 3.1 to describe the set $\mathcal{E}$ in (7) explicitly, see Lemma A.20.
=$\begin{split}\mathcal{E}&=\left\{C\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ C\geq 0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}\text{ satisfies property E}\right\}\\ &=\left\{C\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ C\geq 0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}\text{ satisfies property E}\right\}\end{split}$ (8)
@@
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be injective covariances of equivalent Gaussian measures. Then there exists a sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}w_{i}$ and the following hold
@@
where the leftmost operators are defined on $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$.
Lemma 3.3 shows that the operator in Theorem 3.1(iii) can equivalently described by the operators in items (ii), (iii) and (iv). The operators in Lemma 3.3 can be seen as generalisation of the notion of an operator pencil, which we formally define below.

### v3

@@ 2 Low-rank posterior covariance approximations
Let $\mathcal{H}$ be a separable Hilbert space over $\mathbb{R}$ of dimension $\dim{\mathcal{H}}\leq\infty$, which models the parameter space. Consider the observation model defined by a continuous linear forward model $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ and additive Gaussian observation error
@@
The covariance $\mathcal{C}_{\textup{obs}}\in\mathcal{B}(\mathbb{R}^{n})_{\mathbb{R}}$ of the observation noise $\zeta$ is positive, and from a frequentist nonparametric perspective, $x^{\dagger}\in\mathcal{H}$ is the unknown true data-generating parameter to recover after observing a realisation of $Y$. By the Gaussian assumption on the noise, it follows that for any fixed $x\in\mathcal{H}$, the likelihood of observing $y$ is proportional to $\exp(-\frac{1}{2}\lVert\mathcal{C}_{\textup{obs}}^{-1/2}(y-Gx)\rVert^{2})$. The Hessian of the negative log-likelihood with respect to $x$ is
@@
We adopt the Bayesian perspective to the problem of inferring $x^{\dagger}$ given the observation $y$ of $Y$, by modeling the unknown $x^{\dagger}$ with an $\mathcal{H}$-valued random variable $X$. Its distribution, the prior distribution, is taken to be a Gaussian measure $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$ on $\mathcal{H}$ and we assume that $X$ and $\zeta$ are independent. As the covariance of a Gaussian measure on $\mathcal{H}$, $\mathcal{C}_{\textup{pr}}$ lies in $L_{1}(\mathcal{H})_{\mathbb{R}}$ and $\mathcal{C}_{\textup{pr}}\geq 0$, hence $\mathcal{C}_{\textup{pr}}$ has a unique nonnegative square root $\mathcal{C}_{\textup{pr}}^{1/2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$. In this work, we make the following assumption.
Assumption 2.1.
We assume that the prior distribution $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$ is nondegenerate on $\mathcal{H}$.
Nondegeneracy of $\mu_{\textup{pr}}$ implies that $\operatorname{supp}(\mu_{\textup{pr}})=\mathcal{H}$, see e.g. [4, Definition 3.6.2], and that $\mathcal{C}_{\textup{pr}}>0$ and $\mathcal{C}_{\textup{pr}}^{1/2}>0$, see Lemma A.23. In particular, $\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{\textup{pr}}^{1/2}$ are injective by Lemma A.4. Hence the inverses $\mathcal{C}_{\textup{pr}}^{-1}$ and $\mathcal{C}_{\textup{pr}}^{-1/2}$ are well-defined bijections $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}\rightarrow\mathcal{H}$ and $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}^{1/2}\rightarrow\mathcal{H}$ respectively. They are self-adjoint, c.f. Definition A.18 and Lemma A.22(ii), and if $\dim{\mathcal{H}}=\infty$, then they are unbounded. The Cameron–Martin space of $\mu_{\textup{pr}}$ is the Hilbert space $(\operatorname{ran}{\mathcal{C}}_{\textup{pr}}^{1/2},\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}})$, see e.g. [4, p. 293], where the Cameron–Martin norm of an element $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ is defined by $\lVert h\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}\coloneqq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert$. As $\mathcal{C}_{\textup{pr}}$ is injective and compact, $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}^{1/2}$ is dense in $\mathcal{H}$ and if $\dim{\mathcal{H}}=\infty$, then $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}^{1/2}$ is a proper dense subspace of $\mathcal{H}$.
Given a realisation $y$ of the random variable $Y$ defined by the observation model, the posterior distribution $\mu_{\textup{pos}}=\mu_{\textup{pos}}(y)$ of $X$ given $Y=y$ is the Gaussian measure $\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$, where
@@
see e.g. [32, Example 6.23]. Equation (3c) should be understood to imply the following two facts: $\operatorname{ran}{\mathcal{C}}_{\textup{pos}}\coloneqq\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+H}=\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$, and $\mathcal{C}_{\textup{pr}}^{-1}+H:\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\rightarrow\mathcal{H}$ is the inverse of (3b). While all nondegenerate Gaussians are equivalent in a finite-dimensional setting, this is no longer true in an infinite-dimensional setting, where in fact it holds that nondegenerate Gaussians that are not equivalent must be mutually singular. By [32, Theorem 6.31], $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are in fact equivalent. In particular, $\mu_{\textup{pos}}$ is a nondegenerate measure and the above properties of $\mathcal{C}_{\textup{pr}}$ also hold for $\mathcal{C}_{\textup{pos}}$. We shall construct approximations to $\mu_{\textup{pos}}$ that are equivalent to $\mu_{\textup{pos}}$.
@@
The equations in (3) motivate certain Gaussian approximations of $\mu_{\textup{pos}}$ that, as we shall see, retain equivalence to $\mu_{\textup{pos}}$. By (3b), $\mathcal{C}_{\textup{pos}}$ is an update of $\mathcal{C}_{\textup{pr}}$ by a nonpositive self-adjoint operator $-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}}$. The range of this update is contained in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and the rank of this update is at most $n$ since $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$. For $r\in\mathbb{N}$, this motivates the rank-constrained approximation of $\mathcal{C}_{\textup{pos}}$ by updating $\mathcal{C}_{\textup{pr}}$ using nonpositive self-adjoint operators of the form $-KK^{*}$, for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and $\mathcal{C}_{\textup{pr}}-KK^{*}>0$. That is, we consider
@@
$\mathscr{C}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\right\},\quad r\in\mathbb{N}$ (4)
@@
Since $\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{obs}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}}\in\mathcal{B}_{00,n}(\mathcal{H})_{\mathbb{R}}$, we have $\mathcal{C}_{\textup{pos}}\in\mathscr{C}_{r}$ for all $r\geq r_{0}$ by (3b), where $r_{0}\coloneqq\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{obs}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}}}\right)\leq n$.
Alternatively, we can consider approximations of $\mu_{\textup{pos}}$ by constructing rank-constrained updates of the prior precision $\mathcal{C}_{\textup{pr}}^{-1}$. By (3c), $\mathcal{C}_{\textup{pr}}^{-1}$ is an update of $\mathcal{C}_{\textup{pr}}^{-1}$ by the Hessian $H$, which is self-adjoint, nonnegative, and has rank at most $n$. For $r\in\mathbb{N}$, we can therefore consider the class of approximations of $\mathcal{C}_{\textup{pos}}^{-1}$ of the form $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$, for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. That is, we consider
@@
$\mathscr{P}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\right\},\quad r\in\mathbb{N}.$ (5)
Since $H\in\mathcal{B}_{00,n}(\mathcal{H})_{\mathbb{R}}$, $\mathcal{C}_{\textup{pos}}^{-1}\in\mathscr{P}_{r}$ for all $r\leq r_{0}$ with ${r_{0}}=\operatorname{rank}\left({H}\right)$. The updates $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ in (5) are defined on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, by definition of the sum of unbounded operators, c.f. Section 1.5.
We note that every operator $SS^{*}$ for $S\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a nonnegative, self-adjoint operator with rank at most $r$, and that every nonnegative operator $T\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$ can be written in this way. Therefore, we could write the above approximations as $\mathcal{C}_{\textup{pr}}-T$ or $\mathcal{C}_{\textup{pr}}^{-1}+T$ for nonnegative $T\in\mathcal{B}_{00,r}(\mathcal{H})$, such that $\mathcal{C}_{\textup{pr}}-T$ is positive and maps into $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. However, the set of nonnegative elements of $\mathcal{B}_{00,r}(\mathcal{H})$ is not convex, since rank is not preserved by convex combinations. By replacing $T$ by $SS^{*}$, we avoid formulating an optimisation problem over a nonconvex set. Indeed, $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is not only convex but is also a Banach space.
The classes $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ are generalisations to a possibly infinite-dimensional setting of those considered in [31]. We search for low-rank approximations of the objects in (3b) and (3c), where ‘low-rank’ refers to the fact that we consider approximations in the classes $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$, for $r<n$ respectively. [7, Section 8] contains two examples which can be analysed in the framework described in this section.
@@ 3 Equivalence and Divergences between Gaussian measures
Since our approximation problems are formulated in the context of statistical inverse problems, and since absolute continuity of the posterior with respect to the prior is important for statistical inference, we require our approximate posteriors to be equivalent to $\mu_{\textup{pos}}$. In Section 3.1, we recall the Feldman–Hajek theorem which gives necessary and sufficient conditions for Gaussian measures to be equivalent, and apply this theorem to the setting described in Section 2. Then, in Section 3.2, we consider certain divergences between equivalent Gaussian measures, which we use to measure the approximation quality of low-rank posterior approximations.
Unless otherwise specified, the proofs of the results below are given in Section B.1.
@@ 3.1 Equivalence between Gaussian measures
Let $\mathcal{H}$ be a Hilbert space and $\mu=\mathcal{N}(m_{1},\mathcal{C}_{1})$ and $\nu=\mathcal{N}(m_{2},\mathcal{C}_{2})$ be Gaussian measures on $\mathcal{H}$. Then $\mu$ and $\nu$ are singular or equivalent, and $\mu$ and $\nu$ are equivalent if and only if the following conditions hold:
@@
Remark 3.2 (Cameron–Martin norm equivalence).
Theorem 3.1 states that the Cameron–Martin spaces $\operatorname{ran}{\mathcal{C}_{i}^{1/2}}$, $i=1,2$, of the Gaussian measures $\mu$ and $\nu$ are equal as subspaces if $\mu$ and $\nu$ are equivalent, see also [4, Proposition 2.7.3]. In fact, the two Cameron–Martin spaces $(\operatorname{ran}{\mathcal{C}_{i}^{1/2}},\lVert\cdot\rVert_{\mathcal{C}_{i}^{-1}})$, $i=1,2$, must then have equivalent Cameron–Martin norms as well. This follows from Lemma A.14 applied to the square root of the two covariances. This fact is mentioned without proof in [26, Proposition B.2] and [6, Proposition B.1].
@@
$\mathcal{E}\coloneqq\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{N}(m_{\textup{pos}},\mathcal{C})\sim\mu_{\textup{pos}}\},$ (6)
and more generally, for $m_{1}\in\mathcal{H}$ and $\mathcal{C}_{1}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ with $\mathcal{C}_{1}>0$,
$\mathcal{E}(m_{1},\mathcal{C}_{1})\coloneqq\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{N}(m_{1},\mathcal{C})\sim\mathcal{N}(m_{1},\mathcal{C}_{1})\}.$ (7)
That is, $\mathcal{E}$ contains those covariances $\mathcal{C}$ such that $\mathcal{N}(m_{\textup{pos}},\mathcal{C})$ is equivalent to $\mu_{\textup{pos}}$ and $\mathcal{E}=\mathcal{E}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$. Since $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are equivalent, we have $\mathcal{C}_{\textup{pr}}\in\mathcal{E}$.
@@
By [4, Lemma 6.3.1(ii)] the set of operators that satisfy property E is closed under taking inverses, adjoints and compositions. Furthermore, since $\mathcal{\mu}_{\textup{pos}}\sim\mu_{\textup{pr}}$, $\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}$ satisfies property E. One can now use Theorem 3.1 to describe the set $\mathcal{E}$ in (6) explicitly, see Lemma A.24:
=$\begin{split}\mathcal{E}&=\left\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{C}>0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}\text{ satisfies property E}\right\}\\ &=\left\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{C}>0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}\text{ satisfies property E}\right\}.\end{split}$ (8)
@@
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be injective covariances of equivalent Gaussian measures. Then there exists a sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}w_{i}$ and the following statements hold:
@@
where the domains of the leftmost operators in each statement are dense and in items (i) and (iii) contain $\operatorname{ran}{\mathcal{C}_{1}^{1/2}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}}$.
If $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are as given in Lemma 3.4, then the operator $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is invertible, by Theorem 3.1 and Lemma A.24. Furthermore, the map $\lambda\mapsto\frac{-\lambda}{1+\lambda}$ is a bijection on $(-1,\infty)$. Thus, each of the pairs $(\lambda_{i},w_{i})$, $(\lambda_{i},v_{i})$, $(\frac{-\lambda_{i}}{1+\lambda_{i}},v_{i})$ and $(\frac{-\lambda_{i}}{1+\lambda_{i}},w_{i})$ determines the other three. Hence, Lemma 3.4 shows that the operator in Theorem 3.1(iii) can equivalently be described by the operators in items (ii), (iii) and (iv), which thus all contain the same information. The operators in Lemma 3.4 can be seen as generalisations of the notion of an operator pencil, which we formally define below.